Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2025 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 27     Entry time: 21 Feb 2025, 23:21
Author: Mohammad Vakili 
Group: SPB 
Subject: Shift summary 

SPB/SFX shift summary user experiment 8039

**X-ray delivery**
-15 keV with Hirex monitoring, 0.5 mJ/pulse, 202 pulses/train at 0.56 MHz pulse repetition rate

**Optical laser delivery**
-810 nm laser for illumination

**Achievements/Observations**
- Collected sample data for ca. 2h 15min using a DFFN 
- XTD2 CRLs optimized and realigned
- Offset mirror M1/M2 optimized for 15 keV
- Repositioned the illumination fiber inside IRU chamber
- Germanium blades were installed for the IRU_Slit_4 (top and bottom)

**Issues**
- PP laser alignment could not be pursued as hexapod had to be repositioned first (interaction region was too close to exit cone)

 

 

 

ELOG V3.1.4-7c3fd00